OpenStudy (anonymous):

Sketch a curve y = f(x) that satisfies: f(0) = 3, and dy/dx =-2. What is f(x)?

OpenStudy (shamil98):

You have a differential equation here with an initial value. All you have to do is evaluate the indefinite integral of: \[\int\limits_{}^{} dy = \int\limits_{}^{} -2 ~~dx\] Then evaluate when x = 0, y = 3, to find the Constant. Plug the constant back into the equation for y, thus giving you your particular solution.

OpenStudy (shamil98):

Make sense?..

OpenStudy (shamil98):

\[\frac{ dy }{ dx } = - 2\] \[dy = -2~~dx\] \[\int\limits_{}^{} dy = \int\limits_{}^{} -2~~dx\] and so on.

OpenStudy (anonymous):

we havent done integrations yet so i dont know how to do those

OpenStudy (anonymous):

is it y = - 2x ?

OpenStudy (anonymous):

?

OpenStudy (shamil98):

yes.

OpenStudy (shamil98):

y = -2x + C

OpenStudy (shamil98):

input x = 0, y = 3 and solve for C now. then plug C back in for the particular solution.

OpenStudy (anonymous):

whats c ?

OpenStudy (shamil98):

a constant.

OpenStudy (anonymous):

then c is 3 right ?

OpenStudy (shamil98):

correct. so you have y = -2x + 3

OpenStudy (shamil98):

Just sketch that now, pretty simple stuff.

OpenStudy (anonymous):

oh thanks

OpenStudy (anonymous):

its a straight line but they asked for curve ?

OpenStudy (anonymous):

does that matter ?

OpenStudy (shamil98):

no it doesn't matter.. that is the curve..

OpenStudy (anonymous):

oh ok thanks